OpenStudy (anonymous):

Choose one of the factors of x6 + 1000 x2 x2 – 10 x4 – 10x2 + 100 x4 + 10x2 + 100 I know the answer is x4-10x2+100 but why is that the answer I'm a bit confused

OpenStudy (mertsj):

\[x^6+1000=(x^2)^3+10^3\]

OpenStudy (mertsj):

Factor it using: \[a^3+b^3=(a+b)(a^2-ab+b^2)\]

OpenStudy (anonymous):

x4 + 10x2 + 100